OpenStudy (anonymous):

The Derivative in Graphing and Applications 23. A closed rectangular container with a square base is to have a volume of 2000cm^3. It costs twice as much per square centimeter for the top and bottom as it does for the sides. Find the dimensions of the container of least cost. PLEASE HELPP!!!

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

\[cost \propto C(x)=4hx+\color{brown}4x^2\] \[C(x)={8000\over x}+4x^2\] at least cost: \(C'(x)=0\)

OpenStudy (paxpolaris):

\[8x-{8000 \over x^2}=0\]\[\implies 8x^3=8000\]\[\implies x=10\]\[\implies h={2000 \over 10^2}=20\]